# Satellite Imagery and AI: A New Era in Ocean Conservation, from Research to Deployment and Impact

### v1

@@ 4 Equity considerations
Real time vessel detections from each these models are provided through Skylight, a free maritime intelligence platform. The purposes of these models specifically, and Skylight more broadly, is to help nations protect their marine resources and promote ocean health for future generations. While we believe that the benefits of this technology (both within our platform and as open source repositories) outweigh potential risks, there is a possibility these models could be used for ignoble ends. We do not have a straightforward response that obviates these concerns and we do not take the decision to open source this technology lightly. We chose to open source these models because we believe that both the machine learning research community and the conservation community using the vessel detections should have complete transparency and full access into the underlying model architectures and logic (both for inference and training). In addition, these models are made possible by the existing open geospatial community, especially NASA and the ESA which provides both historical data used for training and near real time data for inference and monitoring.

### v2

@@ 4 Equity considerations
Real time vessel detections from each these models are provided through Skylight, a free maritime intelligence platform. The purposes of these models specifically, and Skylight more broadly, is to help nations protect their marine resources and promote ocean health. While we believe that the benefits of this technology (both within our platform and as open source repositories) outweigh potential risks, there is a possibility these models could be used for nefarious ends. We do not have a straightforward response that obviates these concerns and we do not take the decision to open source lightly. We chose to open source these models because we believe that both the machine learning research community and the conservation community using the vessel detections should have complete transparency and full access into the underlying models and logic. In addition, these models are made possible by the existing open geospatial community, as well as the data providers (NASA, ESA, USGS) which enables both historical analysis and real-time inference.
@@
Allen Institute for AI (Ai2)
All authors are employees at Ai2. Skylight is a product of Ai2.
Computer Vision Annotation team
Ebenezer Aidoo (Ghana Navy), James Curtis Carter (Ghana Navy), Paige Roberts (Skylight)
Defense Innovation Unit
The Defense Innovation Unit contributed funding which supported creation of the VIIRS, S2 and Landsat vessel detection models, and supported the improvements of the S1 model. The S1 model described in this paper was an extension of a previous model created at Ai2 team that was submitted to the xView3 competition.
NASA & LANCE (VIIRS)
We acknowledge the use of data and imagery from NASA’s Land, Atmosphere Near real-time Capability for EOS (LANCE) system (https://earthdata.nasa.gov/lance), part of NASA’s Earth Observing System Data and Information System (EOSDIS). The data products that were used for VIIRS model are described here: https://lance.modaps.eosdis.nasa.gov/viirs/. The following products are used. For SuomiNPP, VNP02DNB_NRT (light), VNP03DNB_NRT (supporting data), VNP02MOD_NRT (gas flares). For NOAA20, VJ102DNB_NRT (light), VJ103DNB_NRT (supporting data), VJ102MOD_NRT (gas flares). In addition, we use the cloud masks created by the University of Wisconsin SSEC (https://www.earthdata.nasa.gov/learn/find-data/near-real-time/viirs-a) [14, 13]. Additional details and code to use these data can be found on the VIIRS GitHub repository [5].
ESA (European Space Agency) and the Copernicus Data Space Ecosystem (S1 and S2)
Data for the S1 and S2 constellations are available at https://dataspace.copernicus.eu/. Additional details of the specific products used for each satellite model are provided below.
  • S1: Level-1 GRD (Ground Range Detected) data were used. Models were trained with both VV and VH polarization modes.
  • S2: Level-1C data (orthorectified top-of-atmosphere reflectance, with sub-pixel multispectral registration). In addition to the L1C data, we also apply cloud detection to suppress false positives due to clouds. To do so we use the s2 cloud detector (s2cloudless) available from PyPI [30]. More details on this algorithm are available in [31].
@@
USGS (Landsat 8-9)
We acknowledge the use of Landsat 8-9 imagery courtesy of the U.S. Geological Survey https://www.usgs.gov/centers/eros/science. The Landsat 8-9 Collection 2 Level-1 scenes data was obtained via the USGS M2M API client. For more details and code, see the Landsat GitHub repository [8].